# Detection of New Galaxy Candidates at 11 in the JADES Field Using JWST NIRCam

### v2

@@ 3 Selection criteria for high-redshift candidate
The candidates should be newly discovered and have not been detected in any previous JWST studies.
@@
In addition to the above criteria, we fitted each SED twice, first restricting EAZY maximum redshift to 7 (low-redshift) and then varying the redshift between 0 20 (high redshift). We estimated the disparity between the best-fit for the low-redshift and high-redshift fit. For this work, we only consider 9, ruling out the low-redshift model at 3 significance (Harikane et al., 2023).
@@ 4 Results and Discussion
Adopting the selection criteria and SED fitting method, we detect seven new galaxy candidates at redshift 11 in the JADES GOODS-S and GOODS-N field. Below we discuss our primary results.
@@ 4.1 Galaxy sample with 11
Utilizing fits generated by EAZY-py and BAGPIPES, we identify seven new candidates within the redshift range of 11 15 across the GOODS-S and GOODS-N fields. Specifically, the five candidates are detected in GOODS-N, with the remaining two candidates detected in the GOODS-S field. Among the five candidates in GOODS-N, three are located in the southeastern sector, while the remaining two are located in its northwestern region. The two targets found in the GOODS-S field are specifically located within its medium region. Figure 1 illustrates the positions of these candidates in the GOODS-N and GOODS-S fields. Among the seven candidates, three had redshifts between 11 and 12, while the remaining four showed best-fit photometric redshifts exceeding 12. Table 1 lists the newly identified candidates along with best-fit parameters from SED fitting.
All candidates are robust detections characterized by a precise high-redshift solution without any significant competing low-redshift alternatives, with 70-100% of their total probability concentrated within z = 1 around the best-fit solution. Similar quality classification criteria have been adopted in a recent JWST UNCOVER high-redshift survey (Atek et al., 2023b), where candidates meeting such criteria have been labeled as high-quality. Our sample targets the faint end of the luminosity distribution, covering a magnitude range of mF277W = 28.5 and 30.4. Previous JADES surveys have documented systems with F277W AB magnitudes reaching as faint as 31.0 (Hainline et al., 2024). Figure 2, 3, 4, 5, 6, 7, and 8 shows the 1.8 1.8 thumbnail images, best-fit SED from BAGPIPES fit, and the posterior redshift distributions. The red points represent the measured photometry. The best-fit favored high-redshift solutions (represented by blue lines) have been shown along with the best-fit low-redshift solutions (represented by grey lines). The SED modeling and the physical properties of the galaxies derived from BAGPIPES fits are described in Section 4.3.
In our dataset, we identified three targets within the lowest redshift range in our sample, 11 z 12 (EAZY determined), as illustrated in Figure 2, 3, and 4: JADES–53.16238-27.87016 (GOODS-S), JADES– 189.17805+62.18788 (GOODS-N), and JADES– 189.09621+62.24877 (GOODS-N). In the observed 11z12 galaxies, none exhibit substantial flux in the F115W images, thus classifying them as F115W dropout candidates. In the next higher redshift range, our survey identified 4 candidate galaxies with photometric redshifts 12z15 (EAZY determined): JADES–53.12089-27.90810 (GOOD-S), JADES–189.26307+62.20648 (GOODS-N), JADES– 189.29834 +62.21367 (GOODS-N), and JADES– 189.10753+62.24305 (GOODS-N). The corresponding thumbnail images are displayed in Figure 5, 6, and 7. JADES– 53.12089 -27.90810 is identified as an F150W dropout, while JADES– 189.26307+62.20648 is categorized as an F115W dropout. The z13 targets, JADES– 189.29834+62.21367 and JADES– 189.10753 +62.24305 are both categorized as F150W dropouts.
@@ 4.2 Size of galaxies
We estimated the half-light radius for F277W filter using the Kron scale factor of 2.5, as adopted by Bradley et al. (2023b).
@@
0.06 and 0.081 (adopting pixel size of 0.03), with a median of 0.063 . Given the F277W PSF full-width at half maximum of 0.088 , and a radius of 0.044 , all the galaxies are spatially resolved. The angular spreads translate to half-light radii ranging between 0.24 kpc to 0.34 kpc, with a median of 0.25 kpc for our galaxy candidates.
These dimensions are in agreement with the rest-UV sizes reported in the GLASS survey, indicating a median half-light radius of 0.45 0.13 kpc for 7z15 galaxies in the F444W band (Yang et al., 2022), as well as with the CEERS survey, which found a median half-light radius of 0.46 kpc for 9z16 galaxies in the F200W band (Finkelstein et al., 2023).
@@ 4.3 Physical Properties through SED fitting
To derive the physical properties of the identified galaxy candidates, we utilized the state-of-the-art Bayesian SED fitting code BAGPIPES (Carnall et al., 2018) to fit the observed SED. BAGPIPES generates complex model galaxy spectra and fits photometric and spectroscopic observations using the MultiNest nested sampling algorithm (Feroz & Hobson, 2008), thereby generating posterior distributions of galaxy properties for each source in the sample. BAGPIPES is capable of modeling galaxies with various star formation histories (SFHs), including delayed-, constant, and bursts (Lower et al., 2020). The stellar population synthesis models adopted for this work are based on BC03 (Bruzual & Charlot, 2003) version 2016. These models assume the Kroupa (2002) initial mass function (IMF) and include nebular line and continuum emission using Cloudy (Chatzikos et al., 2023). For this study, we selected a delayed- SFH (where SFR() exp()), known for its effectiveness in modeling the majority of the stellar population and accurately estimating stellar masses (Ciesla et al., 2017).
@@
We modified the BAGPIPES source code to expand the default redshift range from = 0–10 to = 0–20, ensuring consistency with the EAZY-py results used to select the galaxy sample (Section 2.2). The BAGPIPES SED-fitting was conducted over a wide range of parameter space. We vary within 0.1 and 14 Gyr, stellar masses within = 4 and 13, and log of metallicities (log Z/Z⊙) within 0.005 and 2.5. The logarithm of the ionization parameter (log U), for accounting nebular line and continuum emission, was allowed to vary within and . For this work we adopted Calzetti dust attenuation curve to reduce the number of free parameters in our fits (Calzetti et al., 2000). We let the dust extinction vary within 0 and 4. Additionally, we introduced a multiplicative factor (1 2) to the dust model to address birth-cloud dust attenuation, which typically doubles around H ii regions compared to the general ISM within the galaxy’s first 10 Myr (Bradley et al., 2023b). To account for this, we capped the maximum age of the birth-cloud at 0.01 Gyr (Giménez-Arteaga et al., 2023). The age parameter was allowed to range from 1 Myr to the age of the Universe. Following a similar approach to the EAZY fitting, we conducted two fits of the observed photometry with BAGPIPES. Initially, we constrained the redshift range to 7. Then, we performed another fitting while allowing the redshift to vary between 0.1 and 20. Figure 2, 3, 4, 5, 6, 7, and 8 display the best-fit model SEDs obtained from the BAGPIPES fitting, accompanied by the corresponding values.
@@
The physical parameters extracted from the BAGPIPES fittings are detailed in Table 1. Among these parameters, the inferred intrinsic stellar masses exhibit a range approximately from to 8.75, while the corresponding star formation rates (SFRs) span from around to 5.1 . Additionally, the specific star formation rates (sSFRs), quantified as , vary within the range of approximately to 1.46. Notably, analogous sSFRs have been documented for galaxy candidates at redshifts in Bradley et al. (2023b).
RA | DEC | zEAZY | zBagpipes | mF277 | SFR | Age | Av | ||||
(deg) | (deg) | (AB mag) | (68%) | (Myr) | (pix/arcsec) | ||||||
GOODS-S | |||||||||||
53.16238 | -27.87016 | 11.57 | 11.46 | 28.8 0.1 | 30 | 0.7 | 8.59 | 4.2 | 200 | 0.45 | 2.1 0.1 |
53.12089 | -27.90810 | 12.13 | 12.07 | 28.5 0.1 | 12 | 1.0 | 7.94 | 2.5 | 80 | 0.02 | 2.4 0.2 |
GOODS-N | |||||||||||
189.17805 | 62.18788 | 11.62 | 11.5 | 30.4 0.1 | 13 | 0.9 | 8.25 | 1.8 | 230 | 0.22 | 2.1 0.2 |
189.26307 | 62.20648 | 12.39 | 11.89 | 29.0 0.1 | 25 | 0.9 | 7.75 | 0.9 | 120 | 0.10 | 2.0 0.3 |
189.29834 | 62.21367 | 14.32 | 13.5 | 30.2 0.2 | 15 | 0.9 | 7.93 | 1.3 | 136 | 0.06 | 2.0 0.2 |
189.10753 | 62.24305 | 13.46 | 13.01 | 29.4 0.1 | 10 | 0.9 | 8.62 | 4.2 | 190 | 0.23 | 2.7 0.3 |
189.09621 | 62.24877 | 11.42 | 11.5 | 29.6 0.2 | 26 | 0.9 | 8.75 | 5.1 | 240 | 0.40 | 2.6 0.4 |
@@
Furthermore, the inferred ages for all galaxy candidates within our sample indicate younger stellar populations, spanning a range of approximately 80 to 240 million years, similar to the ages found in earlier photometric detections of galaxies at (Laporte et al., 2021; Bradley et al., 2023b). Regarding dust extinction, the best-fit values of range from 0.02 to 0.45. These values are consistent with those found in galaxies in the NGDEEP survey (Morales et al., 2024), which were also fitted using BAGPIPES and exhibited values between 0.04 and 1.03. These findings collectively contribute to a deeper understanding of the stellar populations and dust properties within the observed galaxy sample.
@@
Noeske et al. (2007) and Elbaz et al. (2007) demonstrated that galaxies at exhibit a tight correlation between star formation rate and stellar mass, with a scatter of 0.2 dex and a rough proportionality characterized by a logarithmic slope of 0.9. This proportionality is similarly observed at using data from the Sloan Digital Sky Survey (Elbaz et al., 2007). However, the normalization at is lower, reflecting the overall decline in cosmic star formation rate density over time. This tight correlation between SFR and stellar mass, known as the main sequence (MS) of star formation, has been extensively documented in the literature (e.g., Daddi et al., 2007; Noeske et al., 2007; Elbaz et al., 2007). This relationship has been observed in high redshifts, with its evolution parametrized up to (e.g., Speagle et al. 2014). However, establishing the MS at high redshifts poses significant challenges due to various systematic uncertainties and selection effects inherent in compiling representative galaxy samples (e.g., Grazian et al., 2015; Förster Schreiber & Wuyts, 2020; Furtak et al., 2021, 2023). In order to probe the poorly understood Universe at , which is beginning to unfold with the advent of the JWST, we investigate the SFR-stellar mass (SFR–) relation within our sample and compare it with previous studies.
In Figure 10, we present the distribution of our galaxy sample in the SFR– plane, derived using BAGPIPES with a 100 Myr SFR time-scale (as described in Section 4.3). We compare our findings with datasets from the NGDEEP survey for - galaxies (Morales et al., 2024), CEERS for - galaxies (Nakajima et al., 2023), earlier JADES surveys for - galaxies (Curti et al., 2024), and the main sequence of star-forming galaxies at (Salmon et al., 2015; Popesso et al., 2023). Our results closely align with those of established studies. Moreover, when fitting the stellar mass versus SFR for our seven galaxy candidates using linear regression, we obtain a slope of , as depicted in Figure 10. Although the slope of our best-fitting line appears less steep compared to those reported in the aforementioned studies, it is important to acknowledge that this discrepancy could stem from biases in our sample selection.
@@ 4.5 Limitations of SED-fitting method and contaminations
We solely utilize SED fitting tool BAGPIPES to derive the physical properties of our high-redshift sample of galaxies. Though the effectiveness and flexibility of BAGPIPES code has been shown in many previous studies, such as Bradley et al. (2023b); Morales et al. (2024), the physical properties of high-redshift galaxies derived from BAGPIPES may be affected by several limitations and contaminations.
Analyzing SEDs through fitting broadband photometry, which mainly captures the rest-frame UV emissions of potential high-redshift galaxies is one of the limitations of our analysis. This wavelength range primarily highlights massive, short-lived stars within a galaxy, which might not represent the majority of its stellar mass. Due to the limited availability of rest-frame optical photometry, SED fitting is expected to underestimate stellar masses. Specifically, Furtak et al. (2021) demonstrated that relying solely on UV photometry can result in an underestimation of stellar masses by up to 0.6 dex. Stellar masses can be accurately derived from the UV continuum only if a galaxy is 10 Myr old, which is the typical lifetime of massive stars. Beyond this age, accurate stellar mass determination requires continuum detection redward of the Balmer break. Furthermore, relying solely on rest-frame UV photometry can create ambiguities between stellar mass, SFR, and age, as noted by Furtak et al. (2023). To confirm if galaxies at possess an evolved stellar population, as suggested by recent simulations (Mason et al., 2023), JWST/NIRCam imaging at wavelengths m should be supplemented with deep imaging from the Mid-Infrared Instrument (MIRI). This approach will enable the probing of the rest-frame optical emission of these very high-redshift galaxies.
@@
The primary source of contamination in broad-band photometry of high-redshift galaxy samples often stems from low-mass Milky Way stars and brown dwarfs. These celestial objects, especially at low temperatures, can exhibit near-infrared (near-IR) colors akin to those of high-redshift galaxies. Previous studies utilizing the HST have indicated that dropout selection techniques can yield contamination levels of up to 40% from low-redshift sources (Bouwens et al., 2011). Hence, it is imperative to consider this possibility in our analyses. Numerous investigations have delved into the identification of these sources within extragalactic surveys (e.g., Finkelstein et al., 2015; Hainline et al., 2020). Brown dwarf candidates have been identified in extragalactic surveys such as GLASS (Nonino et al., 2023). However, the majority of these candidates remain unresolved owing to the relatively large near-infrared camera (NIRCam) point spread function (PSF) of approximately 0.063 for longer wavelengths ( 2m). We assess our results to gauge the potential for low-redshift contaminations.
Following previous studies, such as, those by Bradley et al. (2023b) and Furtak et al. (2023), we extend our analysis by fitting observed SEDs using both EAZY and BAGPIPES. In this process, redshifts were systematically varied within the range of to bias the photometric redshift solutions towards lower values. For all seven galaxies within our sample, the low-redshift fits yield significantly lower significance (i.e., 9) when compared to their corresponding high-redshift solutions. Moreover, it is noteworthy that all galaxies in our sample are well-resolved in the NIRCam longer wavelength channel, with a half-light radius exceeding 0.06 at the F277W band. This observation effectively dismisses the possibility that the galaxies in our sample are low-mass Milky Way stars or brown dwarfs.
Although our selection criteria for high-redshift samples are meticulously crafted to mitigate contamination from lower-redshift interlopers, the potential still exists for some of our candidates to reside at lower redshifts. Spectroscopic confirmation is imperative to accurately determine the redshifts of these candidates.

### v3

@@ 3 Selection criteria for high-redshift candidates
In addition to the above criteria, we fitted each SED twice, first restricting EAZY maximum redshift to 7 (low-redshift) and then varying the redshift between 0 20 (high redshift; see Figure 2–7). We estimated the disparity between the best-fit for the low-redshift and high-redshift fit. For this work, we only consider 9, ruling out the low-redshift model at 3 significance (Harikane et al., 2023).
@@ 4 Results and Discussion
Adopting the selection criteria and SED fitting method, we detect six new galaxy candidates at redshift 11 in the JADES GOODS-S and GOODS-N field. Below we discuss our primary results.
@@ 4.1 Galaxy sample with 11
Utilizing fits generated by EAZY, we identify six new candidates within the redshift range of 11 15 across the GOODS-S and GOODS-N fields. Specifically, the five candidates are detected in GOODS-N, with the remaining two candidates detected in the GOODS-S field. Among the five candidates in GOODS-N, three are located in the southeastern sector, while the remaining two are located in its northwestern region. The two targets found in the GOODS-S field are specifically located within its medium region. Figure 1 illustrates the positions of these candidates in the GOODS-N and GOODS-S fields. Among the six candidates, three had redshifts between 11 and 12, while the remaining four showed best-fit photometric redshifts exceeding 12. Table 1 lists the newly identified candidates along with best-fit parameters from SED fitting.
All candidates are robust detections characterized by a precise high-redshift solution without any significant competing low-redshift alternatives, with 70-100% of their total probability concentrated within z = 1 around the best-fit solution. Similar quality classification criteria have been adopted in a recent JWST UNCOVER high-redshift survey (Atek et al., 2023a), where candidates meeting such criteria have been labeled as high-quality. Our sample targets the faint end of the luminosity distribution, covering a magnitude range of mF277W = 28.5 and 30.4. Previous JADES surveys have documented systems with F277W AB magnitudes reaching as faint as 31.0 (Hainline et al., 2024). Figure 2, 3, 4, 5, 6, and 7 shows the 1.8 1.8 thumbnail images, best-fit SEDs from BAGPIPES and EAZY, and the posterior redshift distributions. The red points represent the measured photometry. The best-fit favored high-redshift solutions (represented by black solid curves) have been shown along with the best-fit low-redshift solutions (represented by grey solid curves). The SED modeling and the physical properties of the galaxies derived from BAGPIPES fits are described in Section 4.3.
In our dataset, we identified three targets within the lowest redshift range in our sample, 11 z 12 (EAZY determined), as illustrated in Figure 2, 3, and 4: JADES–53.16238-27.87016 (GOODS-S), JADES– 189.17805+62.18788 (GOODS-N), and JADES– 189.09621+62.24877 (GOODS-N). In the observed galaxies, none exhibit substantial flux in the F115W images, thus classifying them as F115W dropout candidates. In the next higher redshift range, our survey identified three more candidate galaxies with photometric redshifts (EAZY determined): JADES–189.26307+62.20648 (GOODS-N), JADES– 189.29834 +62.21367 (GOODS-N), and JADES– 189.10753+62.24305 (GOODS-N). The corresponding thumbnail images are displayed in Figure 5, 6, and 7. JADES– 189.26307+62.20648 is categorized as an F115W dropout. The targets, JADES– 189.29834+62.21367 and JADES– 189.10753+62.24305 are both categorized as F150W dropouts.
@@ 4.2 Size of galaxies
We estimated the half-light radius for F277W filter using the Kron scale factor of 2.5, as adopted by Bradley et al. (2023a).
@@
0.06 and 0.081 (adopting pixel size of 0.03), with a median of 0.06 . Given the F277W PSF full-width at half maximum of 0.088, and a radius of 0.044, all the galaxies are spatially resolved. The angular spreads translate to half-light radii ranging between 0.24 kpc to 0.34 kpc, with a median of 0.25 kpc for our galaxy candidates.
These dimensions are in agreement with the rest-UV sizes reported in the GLASS survey, indicating a median half-light radius of
0.45 0.13 kpc for 7z15 galaxies in the F444W band (Yang et al., 2022), as well as with the CEERS survey, which found a median half-light radius of 0.46 kpc for 9z16 galaxies in the F200W band (Finkelstein et al., 2023).
@@ 4.3 Physical Properties through SED fitting
To derive the physical properties of the identified galaxy candidates, we utilized the state-of-the-art Bayesian SED fitting code BAGPIPES (Carnall et al., 2018) to fit the observed SED. BAGPIPES generates complex model galaxy spectra and fits photometric and spectroscopic observations using nested sampling algorithm (e.g., nautilus sampler; Dittmann, 2024), thereby generating posterior distributions of galaxy properties for each source in the sample. BAGPIPES is capable of modeling galaxies with various star formation histories (SFHs), including delayed-, constant, and bursts (Lower et al., 2020). The stellar population synthesis models adopted for this work are based on BC03 (Bruzual and Charlot, 2003) version 2016. These models assume the Kroupa (2002) initial mass function (IMF) and include nebular line and continuum emission using Cloudy (Chatzikos et al., 2023). For this study, we selected a delayed- SFH (where SFR() ) following Morales et al. (2024), known for its effectiveness in modeling the majority of the stellar population and accurately estimating stellar masses (Ciesla et al., 2017).
@@
We modified the BAGPIPES source code to expand the default redshift range from = to = , ensuring consistency with the EAZY-py results used to select the galaxy sample (Section 2.2). The BAGPIPES SED-fitting was conducted over a wide range of parameter space. We vary within 0.1 and 14 Gyr, stellar masses within = 4 and 13, and log of metallicities (log Z/Z⊙) within 0.005 and 2.5. The logarithm of the ionization parameter (log U), for accounting nebular line and continuum emission, was allowed to vary within and . For this work we adopted Calzetti dust attenuation curve to reduce the number of free parameters in our fits (Calzetti et al., 2000). We let the dust extinction vary within 0 and 4. Additionally, we introduced a multiplicative factor (1 2) to the dust model to address birth-cloud dust attenuation, which typically doubles around H ii regions compared to the general ISM within the galaxy’s first 10 Myr (Bradley et al., 2023a). To account for this, we capped the maximum age of the birth-cloud at 0.01 Gyr (Giménez-Arteaga et al., 2023). The age parameter was allowed to range from 1 Myr to the age of the Universe. Following a similar approach to the EAZY fitting, we conducted two fits of the observed photometry with BAGPIPES. Initially, we constrained the redshift range to 7. Then, we performed another fitting while allowing the redshift to vary between 0.1 and 20. Figure 2, 3, 4, 5, 6, and 7 display the best-fit model SEDs obtained from the BAGPIPES fitting, accompanied by the corresponding values.
@@
The physical parameters extracted from the BAGPIPES fittings are detailed in Table 1. Among these parameters, the inferred intrinsic stellar masses exhibit a range approximately from to 8.75, while the corresponding star formation rates (SFRs) span from around to 5.1 . Additionally, the specific star formation rates (sSFRs), quantified as , vary within the range of approximately to 1.46. Notably, similar sSFRs have been found for galaxy candidates at redshifts in Bradley et al. (2023a).
∗ – Redshifts as listed in JADES catalog.
† – Best-fit redshifts from EAZY fit within .
†† – Best-fit redshifts from BAGPIPES fit within .
| RA | DEC | z | z | z | mF277 | SFR (100 Myr) | SFR (10 Myr) | Age | Av | ||||
| (deg) | (deg) | () | (AB) | (68%) | (Myr) | (pix/arcsec) | |||||||
| GOODS-S | |||||||||||||
| 53.16238 | -27.87016 | 3.5 | 11.57 | 11.46 | 28.8 0.1 | 16 | 0.7 | 8.59 | 4.2 | 2.4 | 200 | 0.45 | 2.1 0.1 |
| GOODS-N | |||||||||||||
| 189.17805 | 62.18788 | 2.91 | 11.62 | 11.5 | 30.4 0.1 | 13 | 0.9 | 8.25 | 1.8 | 1.7 | 230 | 0.22 | 2.1 0.2 |
| 189.26307 | 62.20648 | 9.51 | 12.39 | 11.89 | 29.0 0.1 | 25 | 0.9 | 7.75 | 0.9 | 0.85 | 120 | 0.10 | 2.0 0.3 |
| 189.29834 | 62.21367 | 4.93 | 14.32 | 13.5 | 30.2 0.2 | 15 | 0.9 | 7.93 | 1.3 | 1.25 | 136 | 0.06 | 2.0 0.2 |
| 189.10753 | 62.24305 | 3.42 | 13.46 | 13.01 | 29.4 0.1 | 13 | 0.9 | 8.62 | 4.2 | 4.33 | 190 | 0.23 | 2.7 0.3 |
| 189.09621 | 62.24877 | 2.93 | 11.42 | 11.5 | 29.6 0.2 | 26 | 0.9 | 8.75 | 5.1 | 4.86 | 240 | 0.40 | 2.6 0.4 |
@@
Furthermore, the inferred ages for all galaxy candidates within our sample indicate younger stellar populations, spanning a range of approximately 80 to 240 million years, similar to the ages found in earlier photometric detections of galaxies at (Laporte et al., 2021; Bradley et al., 2023a). Regarding dust extinction, the best-fit values of range from 0.02 to 0.45. These values are consistent with those found in galaxies in the NGDEEP survey (Morales et al., 2024), which were also fitted using BAGPIPES and exhibited values between 0.04 and 1.03. These findings collectively contribute to a deeper understanding of the stellar populations and dust properties within the observed galaxy sample.
Recent studies also points to bursty star formation histories in galaxies at compared to their lower-redshift counterparts, with episodic bursts temporarily enhancing the luminosity of low-mass galaxies due to the rapid formation of massive O stars (Looser et al., 2025). At higher redshifts (), several theoretical models have demonstrated that an evolving burstiness can naturally explain the unexpectedly high abundance of bright galaxies detected by JWST in the range (Ferrara et al., 2023). These models suggest that short, intense episodes of star formation—dominate the early buildup of stellar mass in the first few hundred million years after the Big Bang. Therefore, to further assess the impact of star formation history assumptions, we re-derived the stellar mass, redshift, and SFR of our galaxy candidates using a bursty SFH model implemented in BAGPIPES. We implemented the same fitting method as discussed earlier for the delayed- SFH model. Figure 9 presents a comparison between the delayed- and bursty SFH results. The derived stellar masses and redshifts are consistent across both models. However, in 4 out of 6 galaxy candidates, the SFRs inferred from the bursty SFH are significantly higher than those obtained with the delayed- model.
@@
Noeske et al. (2007) and Elbaz et al. (2007) demonstrated that galaxies at exhibit a tight correlation between star formation rate and stellar mass, with a scatter of 0.2 dex and a rough proportionality characterized by a logarithmic slope of 0.9. This proportionality is similarly observed at using data from the Sloan Digital Sky Survey (Elbaz et al., 2007). However, the normalization at is lower, reflecting the overall decline in cosmic star formation rate density over time. This tight correlation between SFR and stellar mass, known as the main sequence (MS) of star formation, has been extensively documented in the literature (e.g., Daddi et al., 2007; Noeske et al., 2007; Elbaz et al., 2007). This relationship has been observed in high redshifts, with its evolution parametrized up to (e.g., Speagle et al. 2014). However, establishing the MS at high redshifts poses significant challenges due to various systematic uncertainties and selection effects inherent in compiling representative galaxy samples (e.g., Grazian et al., 2015; Förster Schreiber and Wuyts, 2020; Furtak et al., 2021, 2023). In order to probe the poorly understood Universe at , which is beginning to unfold with the advent of the JWST, we investigate the SFR-stellar mass (SFR–) relation within our sample and compare it with previous studies.
In Figure 10, we present the distribution of our galaxy sample in the SFR– plane, where the physical properties are derived using the BAGPIPES code. For each galaxy, we compute the SFR averaged over both the most recent 100 Myr and 10 Myr time intervals, allowing us to assess recent star formation activity on different timescales. In the left panel of Figure 10, we compare our 100 Myr-averaged SFR– distribution to that of galaxies identified in the NGDEEP survey spanning redshifts (Morales et al., 2024). Their analysis employs a methodology comparable to ours, using SED fitting to estimate physical properties from deep photometric data, including BAGPIPES code and 100 Myr-averaged SFR. This comparison provides a useful benchmark against another photometrically selected high-redshift sample. Observationally, it has been shown that SFRs derived from dust attenuation corrected H flux are strongly correlated with SFRs averaged over the most recent 10 Myr (Guo et al., 2016; Nakajima et al., 2023). Motivated by this, we compare our 10 Myr-averaged SFR– relation to that of galaxies from the CEERS survey at , for which SFRs were measured using NIRSpec spectroscopy (Nakajima et al., 2023). These spectroscopically confirmed galaxies provide an independent check on our photometric estimates of SFR– distribution over a broad redshift range. In both comparisons, we overlay the star-forming main sequence at as reported by Salmon et al. (2015) and Popesso et al. (2023), shown as reference lines in each panel. Overall, our results are in good agreement with previous studies, supporting the robustness of our derived stellar masses and SFRs across different SFH assumptions and timescales.
@@ 4.5 Limitations of SED-fitting method and contaminations
We solely utilize SED fitting tool BAGPIPES to derive the physical properties of our high-redshift sample of galaxies. Though the effectiveness and flexibility of BAGPIPES code has been shown in many previous studies, such as Bradley et al. (2023a); Morales et al. (2024), the physical properties of high-redshift galaxies derived from BAGPIPES may be affected by several limitations and contaminations.
Analyzing SEDs through fitting broadband photometry, which mainly captures the rest-frame UV emissions of potential high-redshift galaxies is one of the limitations of our analysis. This wavelength range primarily highlights massive, short-lived stars within a galaxy, which might not represent the majority of its stellar mass. Due to the limited availability of rest-frame optical photometry, SED fitting is expected to underestimate stellar masses. Specifically, Furtak et al. (2021) demonstrated that relying solely on UV photometry can result in an underestimation of stellar masses by up to 0.6 dex. Stellar masses can be accurately derived from the UV continuum only if a galaxy is 10 Myr old, which is the typical lifetime of massive stars. Beyond this age, accurate stellar mass determination requires continuum detection redward of the Balmer break. Furthermore, relying solely on rest-frame UV photometry can create ambiguities between stellar mass, SFR, and age, as noted by Furtak et al. (2023). To confirm if galaxies at possess an evolved stellar population, as suggested by recent simulations (Mason et al., 2023), JWST/NIRCam imaging at wavelengths m should be supplemented with future observations targeting the rest-frame optical emission of these very high-redshift galaxies.
@@
The primary source of contamination in broad-band photometry of high-redshift galaxy samples often stems from faint low-redshift galaxies with strong line emission or Balmer breaks. Although our selection criteria for high-redshift samples are meticulously crafted to mitigate contamination from lower-redshift interlopers, the potential still exists for some of our candidates to reside at lower redshifts. Spectroscopic confirmation is imperative to accurately determine the redshifts of these candidates.